OpenStudy (anonymous):

20b - 16= ? please help

OpenStudy (gorv):

what we have to do??

OpenStudy (anonymous):

We have to do what you have to do. :P

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

It is simple, you need to determine \(b\).. :)

OpenStudy (anonymous):

Our first step should be : To add 16 to both the sides, can you do that?

OpenStudy (anonymous):

the question is to factor the experession completely

OpenStudy (anonymous):

I lose.. :( @gorv

OpenStudy (anonymous):

Then who will write the question? you are writing in question box, without writing the question itself? :P

OpenStudy (anonymous):

Okay, it is simple then also: Can you tell me what factor is common between 20 and 16 ?

OpenStudy (anonymous):

gcf or lcf

OpenStudy (anonymous):

Can you think of any number who will divide both 20 and 16 at the same time, yes GCF.

OpenStudy (anonymous):

*which

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Good.. :)

OpenStudy (anonymous):

So, can I write your equation as : \(4 \times (5b) - 4 \times (4) = 0\), Is it right?

OpenStudy (anonymous):

See : \(4 \times (5b)\) will give you again : \(20b\), right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

As, you are seeing \(4\) in both the terms, so you can take \(4\) as common ? Getting?

OpenStudy (anonymous):

is the final answer 4 or 4b?

OpenStudy (anonymous):

See : \[4(5b - 4) = 0\]

OpenStudy (anonymous):

I have factored out \(4\) outside.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now equate brackets to \(0\) and tell me what can be value of \(b\) ?